# Formation of extremal Reissner-Nordström black holes: insights from numerics

### v1

@@ 3.3.3 Massive scalar field
Figure 9: $C^{1}$ gluing with a massive scalar field. Left: Maximum charge $\mathfrak{q}_{\rm max}$ as a function of $\mathfrak{e}M$ for $\Lambda=0$ and $\mathfrak{m}/\mathfrak{e}=0$ (solid blue), $0.01$ (dashed red, lying nearly on top of the blue curve), $0.1$ (dash-dotted green) and $0.5$ (dotted black). The former two curves attain extremality at $\mathfrak{e}M\approx 360$ and $361$, respectively, denoted by stars (the blue star lies mostly beneath the red one). The latter two curves exhibit a kink when the limiting condition $\sup\partial_{U}r$ becomes important. After the kink, the maximum charge rises very slowly with $\mathfrak{e}M$, reaches a maximum, and decreases slowly. The inset zooms in on this phenomenon for the case of $\mathfrak{m}/\mathfrak{e}=0.5$. The peak in the inset is not the kink but is instead part of the near-flat segment. Right: $\inf r$ (top) and $\sup\partial_{U}r$ (bottom) as a function of $\mathfrak{e}M$ for the gluing solution with $\mathfrak{q}=\mathfrak{q}_{\rm max}(\mathfrak{e}M)$, with $\mathfrak{m}/\mathfrak{e}=0.1$.
@@
For the even Ansatz, when $\mathfrak{m}/\mathfrak{e}$ is increased with $\mathfrak{q}$ and $\mathfrak{e}M$ fixed, $\inf r$ decreases so the solution moves closer to the boundary of the restricted parameter space. This implies that even for very small $\mathfrak{m}/\mathfrak{e}$, the lower bound on $\mathfrak{e}M$ to attain some given $\mathfrak{q}$ grows with $\mathfrak{m}/\mathfrak{e}$. For instance, the minimum value of $\mathfrak{e}M$ required to attain extremality grows from about $360$ at $\mathfrak{m}/\mathfrak{e}=0$ to approximately $374$ at $\mathfrak{m}/\mathfrak{e}=0.0374$. In those cases it is still $\inf r$ that sets the bound on $\mathfrak{e}M$. In the $C^{0}$ case the bounds would be identical, as $\inf r$ would not be affected. For higher $\mathfrak{m}/\mathfrak{e}$, the curve of $\mathfrak{q}_{\rm max}$ versus $\mathfrak{e}M$ experiences a kink, much like in the $C^{0}$ case, beyond which the limiting condition $\sup\partial_{U}r<0$ becomes important. Sample curves are provided in Figure 9. The right panels show how $\inf r$ and $\sup\partial_{U}r$ vary for the solution with the largest value of $\mathfrak{q}$ for which gluing can be achieved. This shows how $\mathfrak{q}_{\rm max}$ is determined by $\inf r>0$ before the kink and $\sup\partial_{U}r<0$ after the kink. Beyond the kink, $\mathfrak{q}_{\rm max}$ grows very slowly with $\mathfrak{e}M$, and, if $\mathfrak{m}/\mathfrak{e}\gtrsim 0.03847$, fails to reach extremality. Thus the behaviour of the solutions looks almost qualitatively identical to Figure 5, except that the curves for different $\mathfrak{m}/\mathfrak{e}$ deviate slightly even for small $\mathfrak{q}_{\rm max}$.
@@
Surprisingly, the odd and polynomial Ansätze behave rather differently. When $\mathfrak{m}/\mathfrak{e}$ is increased, the minimal $\mathfrak{e}M$ necessary to obtain extremality at first decreases, as shown in Figure 10, in stark contrast to the case of the even Ansatz (for which the corresponding plot would just increase monotonically, and diverge at $\mathfrak{m}/\mathfrak{e}=0.03847$). This is because, for $\mathfrak{q}=1$ and fixed $\mathfrak{e}M$, $\inf r$ increases with $\mathfrak{m}/\mathfrak{e}$, unlike for the even Ansatz. Hence one can slightly lower $\mathfrak{e}M$ and still form an extremal black hole, i.e. for small $\mathfrak{m}/\mathfrak{e}$, increasing $\mathfrak{m}/\mathfrak{e}$ makes it slightly easier to achieve extremality. This is intriguing, as we know from [Reall:2024njy] that when $\mathfrak{m}/\mathfrak{e}\geq 1$, $\mathfrak{q}=1$ is impossible. So, for large enough $\mathfrak{m}/\mathfrak{e}$ we would expect the minimum value of $\mathfrak{e}M$ to blow up to infinity as $\mathfrak{q}\to 1$, or even earlier, as we see in the case of the even Ansatz. It turns out that the bound of [Reall:2024njy] is still satisfied, although the way that this is manifested is rather different than for the even Ansatz.
@@
Figure 10: Left: Minimum value of $\mathfrak{e}M$ required to reach extremality with the polynomial Ansatz as a function of $\mathfrak{m}/\mathfrak{e}$. On the right of the plot a second solution exists, which merges with the first solution at $\mathfrak{m}/\mathfrak{e}\approx 0.1714$. Right: Same data but for the odd Ansatz. The main plot shows only the segment where $\sup\partial_{U}r$ is the limiting condition. Here, $\mathfrak{e}M$ appears to blow up as $\mathfrak{m}/\mathfrak{e}$ is increased. The inset shows the data for lower mass-to-charge ratios. When $\inf r$ is the limiting condition, the curve decreases monotonically, but then exhibits a sharp kink when $\sup\partial_{U}r$ takes over.
@@
Suppose we increase $\mathfrak{m}/\mathfrak{e}$, while always taking $\mathfrak{e}M$ to be the lowest value that permits extremality, $\mathfrak{q}=1$. During this process, the gluing solution sees $\sup\partial_{U}r$ approach zero, although initially it is still $\inf r$ that determines $\mathfrak{q}_{\rm max}$. Thus, at first $\mathfrak{e}M$ decreases as $\mathfrak{m}/\mathfrak{e}$ is raised. However, when the limiting condition $\sup\partial_{U}r<0$ becomes relevant, then as $\mathfrak{m}/\mathfrak{e}$ is pushed further up, $\mathfrak{e}M$ must be increased to obtain extremality. This is the reason for the sharp kinks in Figure 10 (only the first kink in the case of the first panel). The mass-to-charge ratio cannot be increased indefinitely: beyond some critical value, valid solutions fail to exist. In the case of the odd Ansatz, our results suggest that $\mathfrak{e}M$ blows up to infinity as $\mathfrak{m}/\mathfrak{e}$ increases (see the right plot of Figure 10). We are unable to find valid solutions with $\mathfrak{m}/\mathfrak{e}\gtrsim 0.02193$. The polynomial Ansatz exhibits a more complicated behaviour: immediately after the kink $\mathfrak{e}M$ increases at a variable rate. Then, at a sufficiently large mass-to-charge ratio, a second valid solution exists. These two solutions persist until they coincide near $\mathfrak{m}/\mathfrak{e}\approx 0.1714$. We are unable to find a valid solution, nor even a candidate solution, at higher $\mathfrak{m}/\mathfrak{e}$.
@@
Given the range of behaviours observed with different Ansätze, and the fact that these results depend on the regularity class, it is difficult to draw general conclusions about characteristic gluing with a massive scalar field. For small $\mathfrak{m}/\mathfrak{e}$, the condition determining the minimum value of $\mathfrak{e}M$ required to achieve a given $\mathfrak{q}$ is $\inf r>0$, and this quantity can either decrease or increase as $\mathfrak{m}/\mathfrak{e}$ is varied (with $\mathfrak{e}M$ and $\mathfrak{q}$ fixed). In all cases, when the limiting condition $\sup\partial_{U}r<0$ becomes important, a change in behaviour is induced, although the details of this change depends on the Ansatz. Nevertheless, there is always an upper bound on the value of $\mathfrak{m}/\mathfrak{e}$ for which extremality can be achieved, and this bound is in agreement with [Reall:2024njy].
@@ 3.4 $C^{2}$ Gluing
We now study the $C^{2}$ gluing problem, in which we additionally constrain the second transverse derivative of the scalar field to vanish at $V=1$. Thus, the profile of the scalar field must depend on five parameters. Unfortunately, searching for solutions in a $5$-dimensional space is significantly more complicated than in a $3$-dimensional space, and our numerical methods have proved far less reliable than in the $C^{1}$ case. Moreover, we cannot easily visualise the search space, so identifying suitable initial guesses is difficult. As a result, below we only consider the even Ansatz (and in some instances the modified even Ansatz), for which we have obtained sufficient data to draw some conclusions.
@@
To study how the various branches behave, it is easiest to first try to find as many solutions as possible with fixed parameters. For example, with $\mathfrak{q}=0.2$, $\mathfrak{e}M=300$, $\mathfrak{m}/\mathfrak{e}=0$, $\Lambda M^{2}=0$, we identified three different solutions. As mentioned above, determining solutions is challenging; therefore, there may well exist solutions we have not found. Nevertheless, the ones we uncovered paint a consistent picture, and we believe our conclusions would continue to hold if all solutions were available to us.
@@
Fixing $\mathfrak{m}=0$, the branches obtained from these three solutions behave in the same way as the unique branches of the $C^{0}$ and $C^{1}$ gluing problems. In particular, for each branch, $\sup\partial_{U}r<0$, for any value of $\Lambda M^{2}$. Moreover, as in the $C^{1}$ case, along each branch $\hat{\alpha}$ does not vary significantly with $\mathfrak{q}$ and $\Lambda M^{2}$, so the second panel of Figure 3 gives a good approximation of how the maximum attainable charge depends on $\mathfrak{e}M$ and $\Lambda M^{2}$. (The variance in $\hat{\alpha}$ is roughly similar to that of the $C^{1}$ case.) The minimum value of $\mathfrak{e}M$ needed to attain extremality is different for each branch: approximately $2543$, $2768$ and $3068$ respectively, for $\Lambda=0$. These bounds can be lowered by modifying the Ansatz as in (24). We find an optimal value of $\gamma\approx 0.43$ for each branch. (The optimal values are not exactly identical for each branch, but are very similar; within about half a percent of each other.) This allows us to find a $C^{2}$ gluing solution with $\mathfrak{q}=1$ and $\Lambda=0$ for $\mathfrak{e}M$ as low as $504$.
@@
Representative profiles of the three branches are shown in Figure 11. The profiles chosen are those which achieve extremality with the minimum value of $\mathfrak{e}M$. As in $C^{0}$ and $C^{1}$ gluing, the profiles reach their maximum amplitude in the first half of $\mathcal{C}$, i.e. for $V<1/2$. Note however that the amplitudes of each peak do not necessarily decrease monotonically; that is the case only for the dashed blue branch. The branch which reaches extremality with the smallest value of $\mathfrak{e}M$ is the solid black one. Its maximum amplitude occurs at the second bump. It is interesting that the black and red branches have similar profiles—the signs of each bump are identical, for instance—but reach extremality at very different values of $\mathfrak{e}M$.
Figure 11: Representative profiles for the three branches of solutions for $C^{2}$ gluing with the even Ansatz. The profiles shown are those reaching extremality with the minimum value of $\mathfrak{e}M$ for each branch: $\mathfrak{e}M=2543$ (solid black), $2768$ (dashed blue) and $3068$ (dash-dotted red).
Let us now consider non-zero $\mathfrak{m}$. We find that as $\mathfrak{m}/\mathfrak{e}$ is varied, the three branches change in different ways—even though these branches are all constructed from the same basic profile. The black branch requires larger $\mathfrak{e}M$ to attain extremality with increasing $\mathfrak{m}/\mathfrak{e}$. On the other hand, for the red branch, increasing $\mathfrak{m}/\mathfrak{e}$ reduces $\mathfrak{e}M$, because a higher mass-to-charge ratio increases $\inf r$. This is the same behaviour seen for small $\mathfrak{m}/\mathfrak{e}$ in the polynomial and odd Ansätze in the $C^{1}$ problem. While in both cases, $\sup\partial_{U}r$ gradually increases, it never actually causes issues, because the two branches annihilate when they merge at around $\mathfrak{m}/\mathfrak{e}=0.0102$, well before $\sup\partial_{U}r$ crosses zero. [^7] Annihilation via merging is a well-known phenomenon in bifurcation theory. For instance, consider the polynomial $f(x;s)=x^{2}+s$ depending on some parameter $s$. For $s<0$ there are two real roots. As $s\to 0$-, the two roots converge, and become one multiple root. There are no real roots for $s>0$. This is illustrated in Figure 12. The blue branch does not merge with any other branch. Increasing $\mathfrak{m}/\mathfrak{e}$ raises $\mathfrak{e}M$. Eventually, $\sup\partial_{U}r\to 0$-, at which point a kink appears (see right panel of Figure 12), beyond which the lower bound on $\mathfrak{e}M$ required to reach extremality shoots up and diverges to infinity. The maximum $\mathfrak{m}/\mathfrak{e}$ of the scalar field with which one can achieve $\mathfrak{q}=1$ is approximately $0.01403$.
@@
Figure 12: Minimum value $(\mathfrak{e}M)_{\rm min}$ necessary to reach extremality as a function of $\mathfrak{m}/\mathfrak{e}$, for the even Ansatz in the $C^{2}$ problem. Here $\Lambda=0$. Left: The branches are styled in the same way as in Figure 11. The solid black and dash-dotted red branches merge and annihilate, while the dashed blue branch diverges to infinity. The sudden increase arises due to $\sup\partial_{U}r$ approaching zero. Right: Zoom-in on the kink of the dashed blue branch.

### v2

@@ 3.3.3 Massive scalar field
Figure 10: $C^{1}$ gluing with a massive scalar field. Left: Maximum charge $\mathfrak{q}_{\rm max}$ as a function of $\mathfrak{e}M$ for $\Lambda=0$ and $\mathfrak{m}/\mathfrak{e}=0$ (solid blue), $0.01$ (dashed red, lying nearly on top of the blue curve), $0.07$ (dash-dotted green) and $0.35$ (dotted black). The former two curves attain extremality at $\mathfrak{e}M\approx 720$ and $727$, respectively, denoted by stars. The latter two curves exhibit a kink when the limiting condition $\sup\partial_{U}r$ becomes important. After the kink, the maximum charge rises very slowly with $\mathfrak{e}M$, reaches a maximum, and decreases slowly. The inset zooms in on this phenomenon for the case of $\mathfrak{m}/\mathfrak{e}=0.35$. The peak in the inset is not the kink but is instead part of the near-flat segment. Right: $\inf r$ (top) and $\sup\partial_{U}r$ (bottom) as a function of $\mathfrak{e}M$ for the gluing solution with $\mathfrak{q}=\mathfrak{q}_{\rm max}(\mathfrak{e}M)$, with $\mathfrak{m}/\mathfrak{e}=0.07$.
@@
For the even Ansatz, when $\mathfrak{m}/\mathfrak{e}$ is increased with $\mathfrak{q}$ and $\mathfrak{e}M$ fixed, $\inf r$ decreases so the solution moves closer to the boundary of the restricted parameter space. This implies that even for very small $\mathfrak{m}/\mathfrak{e}$, the lower bound on $\mathfrak{e}M$ to attain some given $\mathfrak{q}$ grows with $\mathfrak{m}/\mathfrak{e}$. For instance, the minimum value of $\mathfrak{e}M$ required to attain extremality grows from about $720$ at $\mathfrak{m}/\mathfrak{e}=0$ to approximately $786$ at $\mathfrak{m}/\mathfrak{e}=0.0265$. In those cases it is still $\inf r$ that sets the bound on $\mathfrak{e}M$. In the $C^{0}$ case the bounds would be identical, as $\inf r$ would not be affected. For higher $\mathfrak{m}/\mathfrak{e}$, the curve of $\mathfrak{q}_{\rm max}$ versus $\mathfrak{e}M$ experiences a kink, much like in the $C^{0}$ case, beyond which the limiting condition $\sup\partial_{U}r<0$ becomes important. Sample curves are provided in Figure 10. The right panels show how $\inf r$ and $\sup\partial_{U}r$ vary for the solution with the largest value of $\mathfrak{q}$ for which gluing can be achieved. This shows how $\mathfrak{q}_{\rm max}$ is determined by $\inf r>0$ before the kink and $\sup\partial_{U}r<0$ after the kink. Beyond the kink, $\mathfrak{q}_{\rm max}$ grows very slowly with $\mathfrak{e}M$, and, if $\mathfrak{m}/\mathfrak{e}\gtrsim 0.02720$, fails to reach extremality. Thus the behaviour of the solutions looks almost qualitatively identical to Figure 5, except that the curves for different $\mathfrak{m}/\mathfrak{e}$ deviate slightly even for small $\mathfrak{q}_{\rm max}$.
@@
Surprisingly, the odd and polynomial Ansätze behave rather differently. When $\mathfrak{m}/\mathfrak{e}$ is increased, the minimal $\mathfrak{e}M$ necessary to obtain extremality at first decreases, as shown in Figure 11, in stark contrast to the case of the even Ansatz (for which the corresponding plot would just increase monotonically, and diverge at $\mathfrak{m}/\mathfrak{e}=0.02720$). This is because, for $\mathfrak{q}=1$ and fixed $\mathfrak{e}M$, $\inf r$ increases with $\mathfrak{m}/\mathfrak{e}$, unlike for the even Ansatz. Hence one can slightly lower $\mathfrak{e}M$ and still form an extremal black hole, i.e. for small $\mathfrak{m}/\mathfrak{e}$, increasing $\mathfrak{m}/\mathfrak{e}$ makes it slightly easier to achieve extremality. This is intriguing, as we know from [17] that when $\mathfrak{m}/\mathfrak{e}\geq 1$, $\mathfrak{q}=1$ is impossible. So, for large enough $\mathfrak{m}/\mathfrak{e}$ we would expect the minimum value of $\mathfrak{e}M$ to blow up to infinity as $\mathfrak{q}\to 1$, or even earlier, as we see in the case of the even Ansatz. It turns out that the bound of [17] is still satisfied, as is shown in Figure 11.
@@
Figure 11: Left: Minimum value of $\mathfrak{e}M$ required to reach extremality with the polynomial Ansatz as a function of $\mathfrak{m}/\mathfrak{e}$. The curve diverges for sufficicntly large $\mathfrak{m}/\mathfrak{e}$. Right: Same data but for the odd Ansatz. The main plot shows only the segment where $\sup\partial_{U}r$ is the limiting condition. Here, $\mathfrak{e}M$ appears to blow up as $\mathfrak{m}/\mathfrak{e}$ is increased. The inset shows the data for lower mass-to-charge ratios. When $\inf r$ is the limiting condition, the curve decreases monotonically, but then exhibits a sharp kink when $\sup\partial_{U}r$ takes over.
@@
Suppose we increase $\mathfrak{m}/\mathfrak{e}$, while always taking $\mathfrak{e}M$ to be the lowest value that permits extremality, $\mathfrak{q}=1$. During this process, the gluing solution sees $\sup\partial_{U}r$ approach zero, although initially it is still $\inf r$ that determines $\mathfrak{q}_{\rm max}$. Thus, at first $\mathfrak{e}M$ decreases as $\mathfrak{m}/\mathfrak{e}$ is raised. However, when the limiting condition $\sup\partial_{U}r<0$ becomes relevant, then as $\mathfrak{m}/\mathfrak{e}$ is pushed further up, $\mathfrak{e}M$ must be increased to obtain extremality. For the odd Ansatz, this leads to the sharp kink visible in the right panel of Figure 11. In contrast to both the even and odd Ansätze, we have not identified any kinks in the curve of the polynomial Ansatz. In all cases, the mass-to-charge ratio cannot be increased indefinitely: beyond some critical value, valid solutions fail to exist, as $(\mathfrak{e}M)_{\rm min}$ blows up to infinity as $\mathfrak{m}/\mathfrak{e}$ increases. For the polynomial Ansatz, this occurs at $\mathfrak{m}/\mathfrak{e}\approx 0.1103$, whereas for the odd Ansatz it occurs at $\mathfrak{m}/\mathfrak{e}\approx 0.01550$.
@@
Given the range of behaviours observed with different Ansätze, and the fact that these results depend on the regularity class, it is difficult to draw general conclusions about characteristic gluing with a massive scalar field. For small $\mathfrak{m}/\mathfrak{e}$, the condition determining the minimum value of $\mathfrak{e}M$ required to achieve a given $\mathfrak{q}$ is $\inf r>0$, and this quantity can either decrease or increase as $\mathfrak{m}/\mathfrak{e}$ is varied (with $\mathfrak{e}M$ and $\mathfrak{q}$ fixed). In all cases, when the limiting condition $\sup\partial_{U}r<0$ becomes important, a change in behaviour is induced, although the details of this change depends on the Ansatz. Nevertheless, there is always an upper bound on the value of $\mathfrak{m}/\mathfrak{e}$ for which extremality can be achieved, and this bound is in agreement with [17].
@@ 3.4 $C^{2}$ Gluing
We now study the $C^{2}$ gluing problem, in which we additionally constrain the second transverse derivative of the scalar field to vanish at $V=1$. Thus, the profile of the scalar field must depend on five parameters. Unfortunately, searching for solutions in a $5$-dimensional space is significantly more complicated than in a $3$-dimensional space, and our numerical methods have proved far less reliable than in the $C^{1}$ case. Moreover, we cannot easily visualise the search space, so identifying suitable initial guesses is difficult. As a result, below we only consider the even Ansatz, for which we have obtained sufficient data to draw some conclusions.
@@
To study how the various branches behave, it is easiest to first try to find as many solutions as possible with fixed parameters. The left panel of Figure 12 shows a number of solutions obtained with $\mathfrak{e}M=600$, $\mathfrak{m}/\mathfrak{e}=0$, $\Lambda M^{2}=0$, and also shows how $\inf r$ varies with $\mathfrak{q}$ for each of them. The picture is rather complicated; branches can cease existing due to $\inf r$ becoming negative, or they can sometimes merge with each other, thereby annihilating themselves. [^9] Annihilation via merging is a well-known phenomenon in bifurcation theory. For instance, consider the polynomial $f(x;s)=x^{2}+s$ depending on some parameter $s$. For $s<0$ there are two real roots. As $s\to 0$-, the two roots converge, and become one multiple root. There are no real roots for $s>0$. Due to the complicated structure, our numerical algorithm is sometimes unstable, and thus there may well exist solutions we have not found. For instance, it seems likely that the branches that disappear without crossing $\inf r=0$ merge with another solution branch which—apart from the dashed green branch in Figure 12—our algorithm has not converged to.
Figure 12: Left: $\inf r$ versus $\mathfrak{q}$ of various solution branches at $\mathfrak{e}M=600$, $\mathfrak{m}/\mathfrak{e}=0$, $\Lambda M^{2}=0$. It is likely that there exist more solutions than presented here. Right: Profiles for $C^{2}$ solutions which form an extremal black hole at $\mathfrak{e}M=4417$ (black) and $\mathfrak{e}M=6198$ (blue), with $\mathfrak{m}=0=\Lambda$.
@@
Fixing $\mathfrak{m}=0$, the branches that do not merge with another seem to behave in the same way as the unique branches of the $C^{0}$ and $C^{1}$ gluing problems. In particular, for each branch, $\sup\partial_{U}r<0$, for any value of $\Lambda M^{2}$. Moreover, as in the $C^{1}$ case, along each branch $\hat{\alpha}$ does not vary significantly with $\mathfrak{q}$ and $\Lambda M^{2}$, so the second panel of Figure 3 gives a good approximation of how the maximum attainable charge depends on $\mathfrak{e}M$ and $\Lambda M^{2}$. (The variance in $\hat{\alpha}$ is roughly similar to that of the $C^{1}$ case.) The minimum value of $\mathfrak{e}M$ needed to attain extremality is different for each branch: approximately $4417$ and $6198$ for the black and blue branches (see Figure 12) respectively, with $\Lambda=0$.
@@
Representative profiles of the two branches we have found that attain extremality are shown in the right panel of Figure 12. The profiles chosen are those which achieve extremality with the minimum value of $\mathfrak{e}M$. As in $C^{0}$ and $C^{1}$ gluing, the profiles reach their maximum amplitude in the first half of $\mathcal{C}$, i.e. for $V<1/2$. Note however that the amplitudes of each peak do not decrease monotonically.
Let us now consider non-zero $\mathfrak{m}$. As $\mathfrak{m}/\mathfrak{e}$ is varied, different branches change in different ways—even though these branches are all constructed from the same basic profile. Two examples are shown in Figure 13. In the left panel, the minimum value of $\mathfrak{e}M$ needed to form an extremal black hole increases with $\mathfrak{m}/\mathfrak{e}$. However, this process stops early, as for $\mathfrak{m}/\mathfrak{e}\gtrsim 0.00144$, we cannot find any candidate solutions (in particular, $\inf r>0$ fails). The right panel shows another example, where the behaviour is more akin to that we observe in the $C^{1}$ problem. For low mass-to-charge ratios of the scalar field, $(\mathfrak{e}M)_{\rm min}$ decreases with $\mathfrak{m}/\mathfrak{e}$, but this relationship turns around sharply when $\sup\partial_{U}r\to 0$-, and $(\mathfrak{e}M)_{\rm min}$ then diverges to infinity. The maximum $\mathfrak{m}/\mathfrak{e}$ of the scalar field with which one can achieve $\mathfrak{q}=1$ is approximately $0.00992$. Much like in the lower regularity problems, it is difficult to draw sweeping conclusions about gluing with a massive scalar field, as different branches can behave in very different ways.
@@
Figure 13: Minimum value $(\mathfrak{e}M)_{\rm min}$ necessary to reach extremality as a function of $\mathfrak{m}/\mathfrak{e}$, for two branches of the even Ansatz in the $C^{2}$ problem. Here $\Lambda=0$. Left: The branch ceases to exist at the point. Beyond, we find $\inf r<0$. Right: Here the branch initially has decreasing $(\mathfrak{e}M)_{\rm min}$ but eventually diverges to infinity. The sudden increase arises due to $\sup\partial_{U}r$ approaching zero.